# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v1

@@ 5.1 Simply supported beam
where $b=L-x_{\mathrm{F}}$, and the non-bold $x$ refers to the position along the beam. This partial model represents the beam in the undamaged condition at a reference temperature. The resulting nominal model is directly incorporated in the physics-based branch of the VAE decoder. The Young’s modulus $E$ and the position of the point load $x_{\mathrm{F}}$ are considered as uncertain latent variables, such that $\bm{z}_{\mathrm{x}}=(E,x_{\mathrm{F}})$. It is assumed that the temperature is an observed domain variable such that $\bm{c}=(T)$, and that the vertical spring log-stiffness $\log k_{\mathrm{v}}$ is taken as a class variable representing damage in the structure, such that $\bm{y}=(\log k_{\mathrm{v}})$. Since the class variable $\bm{y}$ represents damage in the structure it will not be quantitatively measurable. In a realistic scenario, observations of the condition of the support on a qualitative scale (e.g. from 00 representing no damage to $5$ denoting a fully damaged support) might be available. In this example we simplistically consider the ground truth value of $\log k_{\mathrm{v}}$ with some added noise, and assume that measurements of $\bm{y}$ will be available during training but not during testing. The prior distributions over $E$ and $x_{\mathrm{F}}$, as well as the ground truth distributions of the domain and class variables $T$ and $\log k_{\mathrm{v}}$ used to generate the training data are summarized in Table 1. The temperature and vertical spring log-stiffness are not latent variables and are therefore not assigned a prior distribution. To ensure physical consistency and to avoid numerical issues, the Young’s modulus is truncated below a small positive value. The load position $x_{\mathrm{F}}$ is restricted to the range $(0,1)$.
@@
Variable Unit Type Prior distribution Ground truth
$E$ Pa Physical $\mathcal{N}(4.0,1.0)$ $\mathcal{U}(2.5,4.5)$
$x_{\mathrm{F}}$ m Physical $\mathcal{N}(0.5,0.2)$ $\mathcal{U}(0.3,0.7)$
$\log k_{\mathrm{v}}$ N/m Class - $\mathcal{U}(6.0,8.0)$
$T$ Co Domain - $\mathcal{U}(-11.0,5.0)$
@@
After training, the disentanglement between physics-grounded and data-driven components is qualitatively assessed by examining the latent space and samples of the reconstructed response of the beam. A particularly useful tool for assessing the learned representation is to traverse the latent space and the space of reconstructions of the VAE. This can be achieved by generating synthetic data while interpolating over a specified generative factor, and setting the remaining generative factors to a constant value. The VAE is then evaluated on the generated data, yielding samples from the latent space, realizations of the reconstructed input $\hat{\bm{x}}$, and the mean physics-based and data-driven components $\hat{\bm{x}}_{\mathrm{p}}$ and $\hat{\bm{x}}_{\mathrm{d}}$. Each generative factor is linearly interpolated within the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentiles of the corresponding ground truth distribution given in Table 1, while the other variables are held constant at a prescribed value. Additional details on the plotting are provided in Appendix A.
The predicted physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, as well as the combined prediction $\hat{\bm{x}}$ are shown in Figure 5. It can be observed that the data-driven component of the reconstruction $\hat{\bm{x}}_{\mathrm{d}}$ (middle row) is invariant to changes in $E$ and $x_{F}$ contributing only a constant deformed shape to the total predicted response. On the other hand, the physics-based model captures variability in both the physics-based generative factors $\bm{s}_{\mathrm{x}}=(E,x_{\mathrm{F}})$, but also domain and class generative factors $\bm{s}_{\mathrm{c}}=(T)$ and $\bm{s}_{\mathrm{y}}=(\log k_{\mathrm{v}})$. In contrast to the behavior of the unconstrained VAE, presented in Section 2.4, here the model preferentially utilizes the known physics.
@@
Figure 5: Mean prediction and ±$\pm 2\sigma$ uncertainty bounds for the physics-based (top row) and data-driven (middle row) components, and combined prediction (bottom row) while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
To highlight the impact of the GRL, the latent space traversals of the unconstrained model and the model trained adversarially are compared in Figure 6. Without adversarial training, both the domain and class latent variables $\bm{z}_{\mathrm{c}}$, $\bm{z}_{\mathrm{y}}$ encode the variability in the Young’s modulus $E$, providing the data-driven decoder components with the information needed to reconstruct this component of the measured response and resulting in the issues discussed in Section 2.4. In contrast, when $\lambda=1/256$, the GRL regularization yields a posterior distribution over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that is invariant to changes in $E$. This is because the features in $\bm{x}$ that are irrelevant for the reconstruction of the domain and class variables are ignored by the encoder.
@@
Figure 6: Visualizations of the VAE latent space during traversal of the generative factors $E$ and $\log k_{\mathrm{v}}$. Each column corresponds to variation of a single generative factor, and each row shows the marginal approximate posterior distribution of a single latent variable.
@@
The results shown previously indicate that the GRL regularization results in a sparse and parsimonious representation of the physical system, and yields domain and class latent variables that are invariant to changes in the underlying physics. The degree of invariance of the domain and class latent variables to non-domain or class influences can be adapted by tuning $\lambda$. A parameter study is performed in order to assess the impact of the GRL coefficient $\lambda$ on the posterior distribution and the quality of the reconstruction. The model is trained for varying values of $\lambda=\{-1,-1/10,-1/100,-1/1000,0,1/1000,1/100,1/10,1\}$, and the metric described in Section 3.3 is computed for each trained model for a linear regressor. To account for the impact of randomness in the synthetic dataset and the initialization of the neural network parameters, the results for each value of $\lambda$ are averaged over multiple runs. The value of the metric corresponding to each subset $(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\in\bm{z}$ and generative factor, plotted as a function of $\lambda$ are shown in Figure 7.
@@
It can be seen that the sign of $\lambda$ determines the nature of the training procedure, with positive values resulting in adversarial training. For negative values of $\lambda$ the training becomes collaborative, in the sense that the encoder attempts to find approximate posterior distributions over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are jointly informative about their respective modality as well as the response measurements $\bm{x}$. The magnitude of $\lambda$ serves to balance the strength of the adversarial or collaborative training. To aid in the interpretation of the results, the behavior of the model during training is classified into four regimes. When $\lambda$ approaches $-1$ from above, the training is strongly collaborative, and the task of minimizing the error in the reconstruction of $\bm{x}$ is prioritized over the auxiliary tasks of reconstructing $\bm{c}$ and $\bm{y}$. This is reflected by the relatively high scores obtained by the subsets $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ for the generative factors $E$ and $x_{\mathrm{F}}$, indicating that the model has learned an entangled representation. For $\lambda\rightarrow 0^{-}$, the auxiliary tasks are prioritized over the main task, and the amount of information about $E$ and $x_{\mathrm{F}}$ that is encoded in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ is limited. In this regime the behavior can be characterized as weakly collaborative. Conversely, for small positive values of $\lambda$ the training becomes weakly adversarial. In this regime the encoder will seek latent codes over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are uninformative about $\bm{x}$. Further increasing the GRL coefficient such that $\lambda\rightarrow 1$ yields a strongly adversarial model. In this regime, any information that can be used to reconstruct the domain and class variables is heavily weighted against the potential improvement in the reconstruction of $\bm{x}$. In this case the encoder fails to capture the variation in either $\log k_{\mathrm{v}}$ or $T$. The strongly adversarial regime is unlikely to be used in practice and is included only for completeness.
5.2 Damped oscillator
This example demonstrates how the adversarial training prevents the model from compensating for all discrepancies between the physics-based model predictions and measurements. Suppose that a mass-spring-dashpot system undergoes damped harmonic motion, starting from an initial displaced position $x_{0}$, with no external excitation. It is assumed that each experiment is performed under varying temperature $T$, which is taken as the domain variable. The temperature affects the spring stiffness through the relationship $k(T)=k_{\mathrm{ref}}+\alpha_{\mathrm{T}}(T_{\mathrm{ref}}-T)$, with $T_{\mathrm{ref}}=20.0$ Co and $\alpha_{\mathrm{T}}=0.01$. The reference spring stiffness at $T_{\mathrm{ref}}=20.0$ Co is assumed known and equal to $k_{\mathrm{ref}}=1.0$ N/m. The mass $m$ is considered unknown and treated as a physics-grounded latent variable to be inferred from data. The damping coefficient $\zeta$ is taken to define the class of the system. Finally, it is assumed that the observations are subject to an unknown confounding influence in the form of small random perturbations of the initial displacement $x_{0}$. The equation of motion describing the system can be written as:
$m\frac{\mathrm{d}^{2}x(t)}{\mathrm{d}t^{2}}+\zeta\frac{\mathrm{d}x(t)}{\mathrm{d}t}+k(T)x(t)=0$ (5.2)
Each triplet of observations is composed of a noisy displacement time series, and noisy scalar measurements of the damping coefficient $\zeta$ and temperature $T$. The latter are considered as class and domain variables respectively, such that $\bm{c}=(T)$ and $\bm{y}=(\zeta)$. Training and testing datasets are generated by drawing samples from the ground truth distribution and generating the oscillator displacement time-series using the true equation of motion shown in Equation 5.2. The synthetic response measurements are subsequently contaminated with i.i.d. realizations of homoskedastic Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}=0.01$ m. The standard deviation of the measurement uncertainty of the domain and class observables are taken as $\sigma_{\mathrm{c}}=0.01$ and $\sigma_{\mathrm{y}}=0.01$ respectively. A summary of the generative factors, the prior distribution and the ground truth distribution is provided in Table 2. Each of the measured time-series is a vector of $64$ measurements, equally spaced within a time interval $t\in[0,10]$ s.
Variable Unit Type Prior distribution Ground truth
$m$ kg Physical $\mathcal{U}(1.0,2.0)$ $\mathcal{U}(1.2,1.8)$
$\zeta$ Class - $\mathcal{U}(0.0,2.0)$
$T$ Co Domain - $\mathcal{U}(0.0,40.0)$
$x_{0}$ m Unknown - $\mathcal{U}(0.9,1.1)$
$x(t)=\cos\left(\sqrt{\frac{k_{\mathrm{ref}}}{m}}t\right)$ (5.3)
The domain and class latent space dimensions are taken as $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$. The proposed model with no adversarial training ($\lambda=-1$) is trained and evaluated on the synthetic example. The reconstructed response obtained from a traversal of the initial displacement $x_{0}$, shown in 8(a), highlights another issue that occurs when combining physics-based and data-driven components in VAE: Although the variability in $x_{0}$ can not be accounted for by the physics-based model, and there is no information in the domain or class variables regarding the value of $x_{0}$, the lack of regularization results in a model that is free to capture the components of the measured displacement stemming from the variability in the initial displacement $x_{0}$. Although in this case the effect is benign, in more complex physical systems it can result in the model learning unknown confounding influences in a non-interpretable black-box manner. When the GRL regularization is utilized (8(b)), the data-driven encoder is unable to capture the variability in $x_{0}$, effectively depriving the data-driven decoder from the information needed to reconstruct this component of the input measurements.
(a) Samples of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Samples of the reconstructed response with adversarial training ($\lambda=1/128$).
Figure 8: Traversing the space of reconstructions for varying initial displacement $x_{0}$. Physics-based model prediction $\hat{\bm{x}}_{\mathrm{p}}$ (left), data-driven model prediction $\hat{\bm{x}}_{\mathrm{d}}$ (middle), and combined prediction $\hat{\bm{x}}$ (right). With $\lambda=-1.0$ (top) the data-driven components in the VAE are free to capture the variability in the initial position, resulting in underestimation of the uncertainty. For $\lambda=1/128$ (bottom) this is prevented by the adversarial training.
The results shown in Figure 8 demonstrate how the unconstrained VAE will compensate discrepancies between the physics-based model prediction and the measurements caused by unknown confounding influences. The reason why this can be detrimental for the learning task is illustrated in 9(a). The unconstrained VAE prediction accounts for the learned confounding influence of $x_{0}$, resulting in underestimation of the prediction uncertainty. In contrast, when the model is trained with the adversarial objective, the encoder is prevented from learning a representation of $x_{0}$. The uncertainty in the predicted displacement stemming from the variability in $x_{0}$ is more accurately accounted for in the uncertainty bounds of the reconstructed response, shown in 9(b) (right). This uncertainty can also be partly attributed to the fact that the mass-spring-dashpot system does not satisfy the additivity assumption described in Section 2.4.
(a) Samples of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Samples of the reconstructed response with adversarial training ($\lambda=1/128$).
Figure 9: The adversarial training prevents the model from learning the unknown confounding influence of the initial displacement $x_{0}$, resulting in more accurate uncertainty quantification.
5.3 Two-span bridge

### v2

@@ 5.1 Beam case study
where $b=L-x_{\mathrm{F}}$, and the non-bold $x$ refers to the position along the beam. This partial model represents the beam in the undamaged condition at a reference temperature. The nominal model is directly incorporated in the physics-based branch of the VAE decoder. The Young’s modulus $E$ and the position of the point load $x_{\mathrm{F}}$ are considered as uncertain latent variables, such that $\bm{z}_{\mathrm{x}}=(E,x_{\mathrm{F}})$. It is assumed that the temperature is an observed domain variable such that $\bm{c}=(T)$, and that the vertical spring log-stiffness $\log k_{\mathrm{v}}$ is taken as a class variable representing damage in the structure, such that $\bm{y}=(\log k_{\mathrm{v}})$. Since the class variable $\bm{y}$ represents damage in the structure it will not be quantitatively measurable. In a realistic scenario, observations of the condition of the support on a qualitative scale (e.g. from 00 representing no damage to $5$ denoting a fully damaged support) might be available. In this example we simplistically consider $\bm{y}$ as the ground truth value of $\log k_{\mathrm{v}}$ with some added noise. The variable symbols, units, types, as well as the prior distributions over the physics-based latent variables and the ground truth distributions of the generative factors used to generate the training data are summarized in Table 1. We additionally provide a reference value which is used to produce the figures and further discussed below. To ensure physical consistency and to avoid numerical issues, the Young’s modulus is truncated below a small positive value, and the load position $x_{\mathrm{F}}$ is restricted to the range $(0,1)$.
@@
Variable Unit Type Prior distribution Ground truth Reference value
$E$ MPa Physical $\mathcal{N}(4.0,1.0)$ $\mathcal{U}(2.5,4.5)$ 3.0
$x_{\mathrm{F}}$ m Physical $\mathcal{N}(0.5,0.04)$ $\mathcal{U}(0.3,0.7)$ 0.5
$\log k_{\mathrm{v}}$ N/m Class - $\mathcal{U}(6.0,8.0)$ 8.0
$T$ Co Domain - $\mathcal{U}(-11.0,5.0)$ 5.0
@@
After training, the disentanglement between physics-grounded and data-driven components is qualitatively assessed by examining the latent space and samples of the reconstructed response of the beam. A particularly useful tool for assessing the learned representation is to traverse the latent space and the space of reconstructions of the VAE. This can be achieved by generating synthetic data while interpolating over a specified generative factor, and setting the remaining generative factors to a constant reference value, provided in Table 1. The VAE is then evaluated on the generated data, yielding samples from the latent space, realizations of the reconstructed input $\hat{\bm{x}}$, and the mean physics-based and data-driven components $\hat{\bm{x}}_{\mathrm{p}}$ and $\hat{\bm{x}}_{\mathrm{d}}$. Each generative factor is linearly interpolated within the $1^{\mathrm{st}}$ and $99^{\mathrm{th}}$ percentiles of the corresponding ground truth distribution. Additional details on the plotting are provided in Appendix A.
The predicted physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, as well as the combined prediction $\hat{\bm{x}}$ are shown in Figure 5. It can be observed that the data-driven component of the reconstruction $\hat{\bm{x}}_{\mathrm{d}}$ (middle row) is invariant to changes in $E$ and $x_{F}$ contributing only a constant deformed shape to the total predicted response. On the other hand, the physics-based model captures variability in both the physics-based generative factors $\bm{s}_{\mathrm{x}}=(E,x_{\mathrm{F}})$, but also domain and class generative factors $\bm{s}_{\mathrm{c}}=(T)$ and $\bm{s}_{\mathrm{y}}=(\log k_{\mathrm{v}})$. In contrast to the behavior of the unconstrained VAE, presented in Section 2.4, here the model preferentially utilizes the known physics. A key aspect of the adversarial training is the degree to which it allows interaction between the physics-based and data-driven components of the prediction. In this case study, the additional displacement of the beam due to the reduced vertical stiffness of the right-hand side support will also depend on the load position $x_{\mathrm{F}}$. More positive values of $\lambda$ tend to prevent the model from capturing this interaction, whereas more negative values enable it but may result in the data-driven components overriding the known physics.
@@
Figure 5: Mean prediction and ±$\pm 2\sigma$ uncertainty bounds for the physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, and combined prediction $\hat{\bm{x}}$ while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
To further highlight the impact of the GRL, the latent space traversals of the unconstrained model and the model trained adversarially are compared in Figure 6. Without adversarial training, the domain latent variables $\bm{z}_{\mathrm{c}}$ encode the variability in the load position $x_{\mathrm{F}}$, providing the data-driven decoder components with the information needed to reconstruct this component of the measured response and resulting in the issue discussed in Section 2.4. In contrast, when $\lambda=1/256$, the adversarial training results in a posterior distribution over $\bm{z}_{\mathrm{c}}$ that is invariant to changes in $x_{\mathrm{F}}$. The results shown previously indicate that the latent bottleneck architecture and GRL regularization result in a sparse and parsimonious representation of the physical system, and can yield domain and class latent variables that are invariant to changes in the underlying physics. The influence of the GRL hyperparameter $\lambda$ is further investigated in the oscillator case study presented below.
@@
Figure 6: Visualizations of the VAE latent space during traversal of the generative factors $x_{\mathrm{F}}$ and $\log k_{\mathrm{v}}$. Each column corresponds to variation of a single generative factor, and each row shows the marginal approximate posterior distribution of a single latent variable.
5.2 Oscillator case study
This example demonstrates how the adversarial training prevents the model from compensating for all discrepancies between the physics-based model predictions and measurements. Suppose that a mass-spring-dashpot system undergoes damped harmonic motion, starting from an initial displaced position $x_{0}$, with no external excitation. It is assumed that each experiment is performed under varying temperature $T$, which is taken as the domain variable. The temperature affects the spring stiffness through the relationship $k(T)=k_{\mathrm{ref}}+\alpha_{\mathrm{T}}(T_{\mathrm{ref}}-T)$, with $T_{\mathrm{ref}}=20.0$ Co and $\alpha_{\mathrm{T}}=0.01$. The reference spring stiffness at $T_{\mathrm{ref}}=20.0$ Co is assumed known and equal to $k_{\mathrm{ref}}=1.0$ N/m. The mass $m$ is considered unknown and treated as a physics-grounded latent variable to be inferred from data. The viscous damping coefficient $c_{\mathrm{d}}$ is taken to define the class of the system. Finally, it is assumed that the observations are subject to an unknown confounding influence in the form of small random perturbations of the initial displacement $x_{0}$. The equation of motion describing the system can be written as:
$m\frac{\mathrm{d}^{2}x(t)}{\mathrm{d}t^{2}}+c_{\mathrm{d}}\frac{\mathrm{d}x(t)}{\mathrm{d}t}+k(T)x(t)=0$ (5.2)
Each triplet of observations is composed of a noisy displacement time series, and noisy measurements of the viscous damping coefficient $c_{\mathrm{d}}$ and temperature $T$, which are considered as class and domain variables respectively such that $\bm{c}=(T)$ and $\bm{y}=(c_{\mathrm{d}})$. Training and validation datasets are generated by drawing samples from the ground truth distribution and generating the oscillator displacement time-series using the equation of motion shown in Equation 5.2. The synthetic response measurements are subsequently contaminated with i.i.d. realizations of Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}=0.01$ m. The standard deviation of the measurement uncertainty of the domain and class observables are taken as $\sigma_{\mathrm{c}}=0.01$ and $\sigma_{\mathrm{y}}=0.01$ respectively. A summary of the generative factors, the prior distribution and the ground truth distribution is provided in Table 2. Each of the measured time-series is a vector of $64$ measurements, equally spaced within a time interval $t\in[0,10]$ s.
Variable Unit Type Prior distribution Ground truth Reference value
$m$ kg Physical $\mathcal{U}(1.0,2.0)$ $\mathcal{U}(1.2,1.8)$ 1.5
$c_{\mathrm{d}}$ kg / s Class - $\mathcal{U}(0.0,2.0)$ 0.0
$T$ Co Domain - $\mathcal{U}(0.0,40.0)$ 20.0
$x_{0}$ m Unknown - $\mathcal{U}(0.9,1.1)$ 1.0
$x(t)=\cos\left(\sqrt{\frac{k_{\mathrm{ref}}}{m}}t\right)$ (5.3)
The domain and class latent space dimensions are taken as $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$. The proposed model with no adversarial training ($\lambda=-1$) is trained and evaluated on the synthetic example. The reconstructed response obtained from a traversal of the initial displacement $x_{0}$, shown in 7(a), highlights another issue that occurs when combining physics-based and data-driven components in VAE: Although the variability in $x_{0}$ can not be accounted for by the physics-based model, and there is no information in the domain or class variables regarding the value of $x_{0}$, the lack of regularization results in a model that is free to capture the components of the measured displacement stemming from the variability in the initial displacement $x_{0}$. Although in this case the effect is benign, in more complex physical systems it can result in the model learning unknown confounding influences in a non-interpretable black-box manner. When the GRL regularization is utilized (7(b)), the data-driven encoder is unable to capture the variability in $x_{0}$, depriving the data-driven decoder from the information needed to reconstruct this component of the input measurements.
(a) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response with adversarial training ($\lambda=1/128$).
Figure 7: Physics-based model prediction $\hat{\bm{x}}_{\mathrm{p}}$, data-driven model prediction $\hat{\bm{x}}_{\mathrm{d}}$, and combined prediction $\hat{\bm{x}}$ for varying initial displacement $x_{0}$. With $\lambda=-1.0$ (top) the data-driven components in the VAE are free to account for the variability in the initial position. For $\lambda=1/128$ (bottom) the model does not learn this component of the response.
The results shown in Figure 7 demonstrate how the unconstrained VAE will compensate for discrepancies between the physics-based model prediction and the measurements caused by unknown confounding influences. The reason why this can be detrimental for the learning task is illustrated in 8(a). The unconstrained VAE accounts for the learned confounding influence of $x_{0}$, which can lead to underestimation of the uncertainty over the latent variables and predictions. In contrast, when the model is trained with the adversarial objective, the encoder is prevented from learning a representation of $x_{0}$. The uncertainty stemming from the partial knowledge of the physics, including the unknown influence of the viscous damping and the variability in $x_{0}$, is more accurately accounted for in the reconstructed response, as shown in (8(b)). The additional uncertainty can also be attributed to the fact that the mass-spring-dashpot system does not satisfy the additivity assumption described in Section 2.4.
(a) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response with adversarial training ($\lambda=1/128$).
Figure 8: Physics-based model prediction $\hat{\bm{x}}_{\mathrm{p}}$, data-driven model prediction $\hat{\bm{x}}_{\mathrm{d}}$, and combined prediction $\hat{\bm{x}}$ for varying viscous damping coefficient $c_{\mathrm{d}}$. The data-driven decoder components are prevented from fully accounting for the discrepancies between the physics-based model and measurements, resulting in wider uncertainty bounds for the proposed model.
@@ 5.2 Oscillator case study
The trade-off between invariance of the domain and class latent variables to non-domain or class influences and prediction accuracy can be adapted by tuning the GRL hyperparameter $\lambda$. A parameter study is performed in order to assess the impact of different choices for $\lambda$ on the learned representation. The model is trained for varying values of $\lambda=\{-1,-1/10,-1/100,-1/1000,0,1/1000,1/100,1/10,1\}$, and the metric described in Section 3.3 is computed for each trained model using linear regression. The training and testing datasets $D$ and $D^{\prime}$ are composed of $2048$ samples each. To account for the impact of randomness in the synthetic dataset, neural network parameter initialization, and training procedure, the results for each value of $\lambda$ are averaged over multiple runs. The values of the metric for each generative factor and subset of the latent space, as a function of $\lambda$, are shown in Figure 9.
@@
It can be seen that the sign of $\lambda$ determines the nature of the training procedure, with positive values resulting in adversarial training. For negative values of $\lambda$ the training becomes collaborative, in the sense that the encoder attempts to find approximate posterior distributions over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are jointly informative about their respective modality as well as the response measurements $\bm{x}$. The magnitude of $\lambda$ determines the strength of the adversarial or collaborative training. To aid in the interpretation of the results, the behavior of the model is classified into four regimes. When $\lambda$ approaches $-1$ from above, the training is strongly collaborative, and the tasks of minimizing the error in the reconstruction of $\bm{x}$ and the prediction of the domain and class variables $\bm{c}$ and $\bm{y}$ are jointly prioritized. This is reflected by the relatively high scores obtained by the subsets $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ for the generative factors $m$ and $x_{\mathrm{0}}$. For $\lambda\rightarrow 0^{-}$, the auxiliary tasks are prioritized over the main task, and the amount of information about $m$ and $x_{\mathrm{0}}$ that is encoded in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ is limited. In this regime the behavior can be characterized as weakly collaborative. Conversely, for small positive values of $\lambda$ the training becomes weakly adversarial. In this regime the encoder will seek latent codes over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are uninformative about $\bm{x}$. Further increasing the GRL coefficient such that $\lambda\rightarrow 1$ yields a strongly adversarial model, and any information that can be used to reconstruct the domain and class variables is heavily weighted against the potential improvement in the reconstruction of $\bm{x}$. In this case the encoder fails to capture the variation in any of the generative factors.
5.3 Bridge case study